OpenStudy (anonymous):

A ball is thrown vertically upward. After tseconds, its height h (in feet) is given by the function h(t)=96t-16t^2 . After how long will it reach its maximum height? Do not round your answer

OpenStudy (er.mohd.amir):

use diff calculas d/dt(h)=96-32t d2/dt2=-32 equating d/dt(h)=0 give 96-32t=0 t=3 sec

OpenStudy (anonymous):

@Er.Mohd.AMIR thank you